OpenStudy (somegirl1999):

Use pi symbol = 3.14 to estimate the circumference of the circle to the nearest hundredth. circle (1 point) 12 in. 12.56 in. 25.12 in. 50.24 in.2

OpenStudy (anonymous):

Hi :3

OpenStudy (somegirl1999):

Lol Hi

Directrix (directrix):

@SomeGirl1999 Is there a diagram for this problem?

OpenStudy (somegirl1999):

oh yea hold on

Directrix (directrix):

C = 2*pi*r C = 2*3.14*r And, then ?

OpenStudy (somegirl1999):



OpenStudy (somegirl1999):

that is supposed to be a circle lol

Directrix (directrix):

A circle with radius 4?

OpenStudy (somegirl1999):

yea

Directrix (directrix):

C = 2*pi*r C = 2*3.14*r C = 2 * 3.14 * 4 Crank this out and see what you get. Okay?

OpenStudy (somegirl1999):

Thanx for the help i got it

Directrix (directrix):

Glad to help.